Please help me solve this!
Mobile phone companies sometimes package their products to make them more attractive
to customers. If you average 365 minutes per month in talk time, which of these packages is
the better deal?
The Upstart Package includes a free phone and 300 minutes. It will cost 12 cents per minute
for each minute over the plan time. This package has a base rate of $41.95.
The Newstart Package has a phone that costs $10 and has 350 minutes. It will cost 8 cents
per minute for each minute over the plan time. This package has a base rate of $37.95 not including the phone.

Answer by ankor@dixie-net.com(22740)

If you average 365 minutes per month in talk time, which of these packages is
the better deal?
The Upstart Package includes a free phone and 300 minutes.
It will cost 12 cents per minute for each minute over the plan time.
This package has a base rate of $41.95.
.12(365-300) + 41.95 =
.12(65) + 41.95 =
7.80 + 41.95 = $49.75
:
The Newstart Package has a phone that costs $10 and has 350 minutes.
It will cost 8 cents per minute for each minute over the plan time.
This package has a base rate of $37.95 not including the phone.
.08(365-350) + 10 + 37.95 =
.08(15) + 47.95 =
1.20 + 47.95 = $49.15